### 22 Feb 2024 v3

# Testing theories of the glass transition with the same liquid, but many kinetic rules

Abstract
We study the glass transition by exploring a broad class of kinetic rules that can significantly modify the normal dynamics of super-cooled liquids, while maintaining thermal equilibrium. Beyond the usual dynamics of liquids, this class includes dynamics in which a fraction $(1-f_{R})$ of the particles can perform pairwise exchange or ‘swap moves’, while a fraction $f_{P}$ of the particles can only move along restricted directions. We find that (i) the location of the glass transition varies greatly but smoothly as $f_{P}$ and $f_{R}$ change and (ii) it is governed by a linear combination of $f_{P}$ and $f_{R}$. (iii) Dynamical heterogeneities (DH) are not governed by the static structure of the material, their magnitude correlates instead with the relaxation time. (iv) We show that a recent theory for temporal growth of DH based on thermal avalanches holds quantitatively throughout the $(f_{R},f_{P})$ diagram. These observations are negative items for some existing theories of the glass transition, particularly those reliant on growing thermodynamic order or locally favored structure, and open new avenues to test other approaches, as we illustrate.
Understanding why liquids glass formers cease to flow near their glass transition $T_{g}$ remains a challenge. At that point, the relaxation time $\tau_{\alpha}$ beyond which stress relaxes is of order of minutes, which is fifteen decades larger than at high temperatures. From $\tau_{\alpha}$, the activation energy $E_{a}$ can be defined as $\tau_{\alpha}=t_{0}\exp(E_{a}/T)$, where $t_{0}$ is a microscopic time scale and $T$ is the temperature (in the units of the Boltzmann constant). In liquids called fragile, $E_{a}$ can increase five-fold or more under cooling Anderson (1995); Ediger et al. (1996); Debenedetti and Stillinger (2001); Berthier and Biroli (2011). As the dynamics slows down, it also becomes more and more heterogeneous, corresponding to a growing length scale $\xi$  Kob et al. (1997); Yamamoto and Onuki (1998); Dalle-Ferrier et al. (2007); Karmakar et al. (2014). Contrasting theories seek to explain these two facts. In the first class of views, including Adam-Gibbs Adam and Gibbs (1965) and Random First Order Theory (RFOT) Kirkpatrick et al. (1989); Lubchenko and Wolynes (2007); Biroli and Bouchaud (2012), the increase of activation energy stems from the emergence of some order on a growing length $\xi$, that must be destroyed by cooperative motion on that scale to relax the material. $E_{a}$ can then be expressed in terms of purely thermodynamic quantities, independently of the kinetic rules governing the dynamics. Some real space approaches associate such a growing order to locally favored structures Patrick Royall et al. (2008); Kawasaki et al. (2007). A second viewpoint seeks to capture the mechanism of dynamical facilitation, whereby the relaxation of a given region speeds up the relaxation of regions nearby. Kinetically constrained models Ritort and Sollich (2003); Berthier et al. (2011), such as the East model, capture this effect and suggest a scenario Garrahan and Chandler (2002); Garrahan et al. (2007); Hedges et al. (2009) in which thermodynamics plays almost no role, but dynamics is heterogeneous and the growth of activation energy stems from non-local rearrangements taking place over $\xi$. At odd with these two views, a third approach that includes free volume Turnbull and Cohen (1961) or elastic Dyre (2006); Kapteijns et al. (2021); Lerbinger et al. (2022); Ciamarra et al. (2023) models, assumes that the activation energy is not controlled by a growing length scale. Instead, it is governed by the energy cost of elementary rearrangements of a few particles jumping over a barrier. The elastic coupling between rearrangements Lemaître (2014); Chowdhury et al. (2016); Tong et al. (2020); Wu et al. (2015); Maier et al. (2017); Steffen et al. (2022); Flenner and Szamel (2015); Klochko et al. (2022); Chacko et al. (2021) leads to a correlated dynamics  Ozawa and Biroli (2023) that can be described in terms of avalanches of activated events Tahaei et al. (2023).
Figure 1: Diagram indicating the number of degrees of freedom for spatial dimension $d=2$ (a) and $d=3$ (c) as a function of the fraction $f_{R}$ of particles that cannot swap and the fraction $f_{P}$ of particles whose translation motion is restricted. Panel (b) shows the particles positions at various time points within some time interval for $d=2$. Particles that can move freely in all direction are shown in red, and particles that are restricted to linear motion appear in blue. For $d=3$, this constraint is more gentle, as particles can still move on planes as sketched in (c).
Molecular dynamics simulations of models of super-cooled liquids have been extremely informative to characterize the glass transition Kob (1999), yet distinct views on this phenomenon have been hard to definitely contrast Biroli and Bouchaud (2012). Our present goal is to show that for some popular models of liquids, a very broad class of kinetic rules can be considered, which can continuously (and very significantly) speed up or slow-down the normal dynamics, while preserving thermal equilibrium. Although these rules would be hard to implement in actual experiments, they are equivalent to dynamics with purely local rules, and as such theories of the glass transition should apply to them. This approach thus opens an avenue to test more stringently theories of glassy dynamics. Specifically, our work builds on ‘swap’ Monte Carlo algorithms. In these algorithms, pairs of particles can exchange positions, in addition to their usual translation moves Briano and Glandt (1984); Grigera and Parisi (2001); Fernández et al. (2007); Gutiérrez et al. (2015); Ninarello et al. (2017). For continuously polydisperse systems, these algorithms can speed up the dynamics by 15 orders of magnitude or more Ninarello et al. (2017), and can change the glass transition temperature $T_{g}$ by up to a factor of two. It allows one to explore glasses with a stability similar to that reached in experiments. Our central result is to introduce a family of kinetic rules, where a fraction $f_{R}$ of the particles cannot swap, and a fraction $f_{P}$ of the particles can only move along randomly-chosen hyperplanes. We provide systematic measurements of the dynamics in the $(f_{R},f_{P})$ diagram in two and three dimensions, that includes the normal dynamics $(1,0)$ as well as swap $(0,0)$. Overall, our observations are negative items for theories based on a growing thermodynamic order. We discuss how devoted studies could be used in these models to test alternative views of the glass transition. Changing continuously the kinetic rules of liquids: Swap moves lead to a considerable speed up of the dynamics Briano and Glandt (1984); Gutiérrez et al. (2015); Ninarello et al. (2017). Importantly, despite its apparent non-local character, swap dynamics can be conceived as a purely local dynamics. Following Brito et al. (2018); Briano and Glandt (1984); Hagh et al. (2022), swap is equivalent to considering identical particles endowed with an additional ‘breathing’ degree of freedom, allowing them to change their size according to some chemical potential $\mu(R)$. Indeed, letting pairs of particles exchange is equivalent, in the thermodynamic limit, to letting individual particles exchange with a bath of particles of all possible sizes $R$. $\mu(R)$ is then chosen to obtain the desired polydispersity, which is continuous for continuously poly-disperse particles Brito et al. (2018); Briano and Glandt (1984). Adding such a degree of freedom per particle dramatically softens the energy landscape Szamel (2019); Ikeda et al. (2019); Brito et al. (2018); Wyart and Cates (2017) while preserving thermodynamic and structural properties. It also affects the dynamics: following the center of the particles, and considering that a swap move corresponds to a change of the size of two particles but not of their position, leads to the following observation. Dynamical correlations grow under cooling as for the normal dynamics, but the correlation length starts growing at a much smaller temperature Ninarello et al. (2017). To study more systematically such effects and their consequences, following Gopinath et al. (2022) we vary the parameter $f_{R}\in[0,1]$, characterizing the number of particles that cannot swap. Unlike Gopinath et al. (2022), we do not consider that particles that swap positions perform a jump, as it leads to very different dynamics. Following this logic, we propose to add even more kinetic constraints by restricting the motion of a fraction $f_{P}\in[0,1]$ of the particles, as illustrated in Fig.1. Each such particle is forbidden to move along one random direction associated to it, and for an infinite system they would be each confined along a random hyperplane. Overall, the number of degrees of freedom for a system of $N$ particles is then $N(d-f_{P}+1-f_{R})$. Note that for the periodic boundary condition we consider below, this dynamics is ergodic as such hyperplanes visit the neighborhood of any point with probability one, if their orientation is randomly chosen. For dynamics that satisfy detailed balance like ours, it ensures that thermal equilibrium will eventually be reached. Thus structural and thermodynamic properties are only governed by $\phi$, independently of the choices of kinetic rules embodied in $(f_{R},f_{P})$. Note that other procedures were proposed to reduce the number of degrees of freedom, such as pinning particles starting from an equilibrated system as proposed e.g. in Cammarota and Biroli (2012). Yet in that case ergodicity is obviously broken once pinned particles are chosen, and the dynamical properties of the system are not translation-invariant anymore. In the present work, we specifically perform Monte Carlo simulations of systems with $N$ continuously polydisperse hard spheres particles of packing fraction $\phi$ in a regular box of linear size $L$, with periodic boundary conditions. For hard particles, instead of temperature, $\phi$ is the good controlled parameter. The choice of polydispersity together with other numerical details are shown in the Appendix A. Figure 1-(a,c) shows a schematic diagram of the different dynamics we explore in $d=2$ and $d=3$ respectively, and indicate in color the associated number of degrees of freedom. Figure1-(b) illustrates an example of the particles trajectories at a short time for $d=2$, revealing which particles are restricted to linear motion, and which ones are not.
Figure 2: $\tau_{\alpha}/\tau_{\alpha}^{*}$ as function of $\phi$ for different values of $f_{R}$ and $f_{P}$ for $d=2$ (a) and $d=3$ (e). From these curves, $\phi_{G}$ as a function of $f_{R}$ and $f_{P}$ is extracted, as indicated in color for $d=2$ (b) and $d=3$ (f). The discs mark the simulation data, from which extrapolations are made. (c,g) show $\phi_{G}$ as a function of the number of degrees of freedom per particle $d+1-f_{R}-f_{P}$ and (d,h) show that $\phi_{G}$ can be collapsed as a function of an effective number of constraints $x=f_{P}+C(d)f_{P}$, with $C(2)=0.45$, and $C(3)=0.70$
Dependence of the glass transition packing fraction $\phi_{G}$ with kinetic rules: For $d=3$, our choice of polydispersity and Monte-Carlo algorithm follows closely previous works Berthier et al. (2016); Ninarello et al. (2017), where it was shown that swap can speed up the dynamics by 15 decades or more. Here we observe a giant speed up of swap in our $d=2$ system as well, as documented in Appendix B. We consider the dynamics on the full phase diagram $(f_{R},f_{P})$. Fig.2-(a,e) show $\tau_{\alpha}$ vs $\phi$ for such dynamics as the parameters $(f_{R},f_{P})$ are varied, both for $d=2$ and $d=3$. Fig.2-(b,f) represents the corresponding value of $\phi_{G}$ in color, as extrapolated from the different measurements made as indicated by circles. The most remarkable results are that (i) $\phi_{G}$ varies very significantly, and continuously throughout the phase diagram. In particular, there is no evidence for a region surrounding the normal dynamics where kinetic constraints would not matter and $\phi_{G}$ would plateau. The normal dynamics can be made continuously faster or slower. (ii) Observing these two diagrams, it is apparent that most of the variation of $\phi_{G}$ is captured by a linear combination of $f_{R}$ and $f_{P}$. This hypothesis is tested in Fig.2-(d,h), where it is shown that $\phi_{G}(f_{R},f_{P})=\phi_{G}(x)$, where $x$ can be thought of as an effective number of constraints $x=f_{P}+C(d)f_{R}$, where the coefficient $C(d)\leqslant 1$ characterize the relative effect of breathing degrees of freedom v.s. translational ones. (iii) qualitative observations are independent of the spatial dimension $d$.
Figure 3: (a) Maximum $\chi_{4}^{*}$ of the dynamic susceptibility $\chi_{4}(t)$ for $d=3$ as a function of $\phi$ as $f_{R}$ as is varied indicated in legend, and $f_{P}=0$. (b) $*\chi_{4}^{*}/\chi_{4}{{}^{*}}(\phi=0.4)$ as a function of $\phi/\phi_{G}$. Comparing (a) and (b) reveals that $\phi/\phi_{G}$ is a much better predictor of DH than $\phi$. (c) Coarsening length $l_{c}$ for $d=2$ v.s. $\ln(t)$ as $\phi$ is varied, for different values of $f_{R}$ indicated in legend (d) for $f_{P}=0$ and $t\leq\tau_{\alpha}/3$. (d) $l_{c}$ collapses when axis are rescaled as predicted by Eq.1. The full black line is the theoretical prediction $l_{c}=a\sqrt{\chi_{4}^{*}}[\ln(\tau_{\alpha}/t)]^{-1}$ with $a=9.5$.
Dynamical heterogeneities are not governed by structural properties: Various theoretical approaches propose that dynamical heterogeneities are controlled by equilibrated structural properties, such as the extension of locally-favored structure Patrick Royall et al. (2008); Kawasaki et al. (2007) or a point to set length Lubchenko and Wolynes (2007); Biroli and Bouchaud (2012) entering RFOT (although additional effects such as facilitation can be added to this theory to increase further dynamical correlations, see e.g. Biroli and Bouchaud (2023) for a recent discussion). In our set-up, these properties depend only on $\phi$, and not on the values of parameters $f_{R},f_{P}$. Here instead we find that for the relaxation times we can probe and the system we consider, the dynamical susceptibility $\chi_{4}$ characterizing the magnitude of dynamical heterogeneities are not governed by $\phi$ only, but instead depend strongly on $f_{R},f_{P}$ as illustrated in Fig.3(a). Much less variation in $\chi_{4}$ is found when plotted as a function of $\phi/\phi_{G}$ as shown in Fig.3(b). These results indicate that heterogeneities are not controlled by a length scale that would directly appear in the structure. Indeed, dynamical correlations are small for the swap dynamics at significant packing fraction for which the normal dynamics is already very correlated Ninarello et al. (2017). Thus, although locally favored structure may affect dynamical heterogeneities in specific systems (such as two-dimensional systems of discs that can display hexatic order Kawasaki et al. (2007)), it does not appear to be the case for continuously polydisperse particles, at least in the range of time scale that can be probed numerically. The temporal evolution of DH follows that of thermal avalanches: Starting from a reference equilibrated configuration, coloring rearranging regions reveals a coarsening phenomenon Chandler and Garrahan (2010); Scalliet et al. (2022) and a growing length scale $l_{c}(t)$, as shown in Fig.3. Quantitatively we define $l_{c}(t)=\langle l^{3}\rangle_{t}/\langle l^{2}\rangle_{t}$ where the average $\langle\rangle_{t}$ is made on all clusters of connected relaxed particles at time $t$, and $l$ is the square root of the number of particles involved in a cluster. In Tahaei et al. (2023), a description of this coarsening based on facilitation via thermal avalanches of local activated rearrangements was proposed. It was refined and also argued to apply to creep flows in elastic manifolds in de Geus et al. (2024), and implies for $t\ll\tau_{\alpha}$ and $l_{c}$ much smaller than the system size:
$l_{c}\sim\xi\left[\ln(\tau_{\alpha}/t)\right]^{-1/(\tilde{\sigma}\tilde{d}_{f})}$ (1)
where $\xi$ is the correlation length, $\tilde{\sigma}$ some exponent and $\tilde{d}_{f}$ is the fractal dimension of the clusters. In this view, $\xi^{\tilde{d}_{f}}\sim\chi_{4}^{*}$ which characterizes the typical cluster volume. For $d=2$, $\tilde{d}_{f}\approx 2$ and $\tilde{\sigma}\approx 1/2$ Tahaei et al. (2023), leading to $l_{c}\sim\sqrt{\chi_{4}^{*}}\left[\ln(\tau_{\alpha}/t)\right]^{-1}$. As shown in in Fig.3(d), this prediction works remarkably well, independently of $f_{R}$. It supports that thermal avalanches cause DH independently of the kinetic rules chosen. Discussion: Swap Monte Carlo algorithms can be restricted to local moves with no significant effects on the dynamics Ninarello et al. (2017). More fundamentally, they are equivalent to a purely local kinetic rule where particles can adapt their radii Briano and Glandt (1984); Brito et al. (2018). For theories based on a growing order on some length $\xi_{coop}$ such as RFOT, or approaches based on locally favored structures, barriers are cooperative and can be expressed in terms of thermodynamic quantities alone. They should be present for any local kinetic rules Wyart and Cates (2017), including those studied here. Thus in these approaches, the core mechanism slowing down the dynamics near the glass transition should not depend on the choice of $(f_{R},f_{P})$. Authors in Berthier et al. (2019) acknowledged that swap and normal dynamics should asymptotically relax at the same pace according to RFOT, but countered that pre-asymptotic corrections (not currently described within this theory) may cause the observed difference. The main difficulty with this view is that the dynamics is very different as the parameters $(f_{R},f_{P})$ change: as shown in Fig.5, these dynamics do not become equivalent even after a slowing down of 15 decades accessible experimentally. To have predictive power, RFOT or theories based on thermodynamic quantities should specify a value for the parameters $(f_{R},f_{P})$ for which they apply. However, currently they don’t. Our observation that $\phi_{G}$ continuously depends on $(f_{R},f_{P})$, and does not plateau to some constant value around the normal dynamics $(1,0)$, shows that the normal dynamics is just one among many other dynamics. This point underlines the lack of predictive power of RFOT or related theories - at least for the systems of continuously polydisperse particles studied here. By contrast, for theories based on kinetic constraints or on local barriers (known to depend on kinetic rules Brito et al. (2018)), the fact that $\phi_{G}$ should continuously vary with the amount of constraints is evident. The normal dynamics is slower simply because it is a kinetically constrained version of swap dynamics. Conclusion: RFOT is a mean-field theory of the glass transition, which has shown undeniable successes. It is exact in infinite dimension Charbonneau et al. (2014); Maimbourg et al. (2016), correctly captures aspects of the thermodynamics of super-cooled liquids Biroli et al. (2008), and presents a dynamical transition Kurchan and Laloux (1996); Biroli et al. (2006); Lubchenko and Wolynes (2007) akin to mode coupling theory, that describes some aspects of liquid dynamics at intermediate temperatures Götze (2008). Yet our results support that its description of activation near the glass transition does not apply for the continuously polydisperse systems studied here. Although our conclusions are restricted to these specific systems where swap is so performant, these models are known to capture the key facts associated with the glass transition Scalliet et al. (2022). The model we introduced, with its very large variation of dynamics with different kinetic rules, offers new opportunities to test different theories of the glass transition, extending previous observations that were only considered for a single kinetic rule. As an illustration, we showed that the time evolution of dynamical heterogeneities follow a description of facilitation based on thermal avalanches of local rearrangements Tahaei et al. (2023) independently of the kinetic rules chosen, emphasizing the robustness of this theory. Other debated issues that this model would help resolve include what observable predicts best the regions which will flow first, the so-called dynamical propensity, for which many candidates were proposed Widmer-Cooper et al. (2004); Lerbinger et al. (2022); Alkemade et al. (2022); Schoenholz et al. (2017). Likewise, numerical tests put forward to test theories of the slowing down of the dynamics (i.e. how the activation energy $E_{a}$ depends on temperature or packing fraction) can now be made much more stringent. The list includes kinetically constrained or lattice gas models Toninelli et al. (2006), that can be tested by analyzing irreversible events Keys et al. (2011). Furthermore, the notion that glassy dynamics correspond to local rearrangements was supported by the measurement of the density of state of local barriers Ciamarra et al. (2023); and these barriers were argued to be governed alternatively by global Dyre (2006) or local Kapteijns et al. (2021) elasticity, or by the amplitude of vibrational motion Puosi and Leporini (2012). Varying $(f_{R},f_{P})$ in these measurements will indicate which viewpoint is most likely correct. Overall, we have added an axis to the glass transition problem by varying continuously kinetic rules, affecting strongly observations and giving a new handle to decide which theory of the glass transition actually applies.
Acknowledgements.
We thank the Simons collaboration for discussions, in particular L. Berthier, G. Biroli, M. Ozawa and C. Scalliet. We thank J.P. Bouchaud, T. deGeus, W. Ji, M. Muller, M. Pica Ciamarra, M. Popovic, A. Tahei and A. Rosso for discussions, and J. Kurchan for exchanges at the beginning of this project. MW acknowledges sup- port from the Simons Foundation Grant (No. 454953 Matthieu Wyart) and from the SNSF under Grant No. 200021-165509. C.B. and C.G. thank the Brazilian agency CAPES and CNPq for the financial support..
Appendix A Numerical Details
We use for system sizes $N_{d=2}=484$ and $N_{d=3}=512$, which allow to perform extensive simulations as $\phi$, $f_{P}$ and $f_{R}$ are varied, while having small finite-size effects for the dynamical range considered Berthier et al. (2012). Our Monte Carlo algorithm follows previous choices Ninarello et al. (2017): it involves displacements and swap moves, where the latter are attempted with a probability of $20\%$. The magnitude $\delta l$ of translation moves is chosen such that the acceptance ratio is $75\%$. To achieve rapidly equilibration for any choice of $(\phi,f_{R},f_{P})$, we first use our fastest Monte Carlo, corresponding to $(\phi,f_{R}=0,f_{P}=0)$. Then our algorithm is run with the desired values of $(f_{R},f_{P})$ for $10^{9}$ Monte Carlo steps. To check that equilibration was reached, we compare relevant observables (such as correlation functions) in the first and second half of the run, and test for consistency. The magnitude of the effect of the SWAP in speeding up the dynamics is affected by the width of the particle radii distribution. The degree of polydispersity is quantified as in Ninarello et al. (2017) by some quantity $\delta=\sqrt{\langle\sigma^{2}\rangle-\langle\sigma\rangle^{2}}/\langle\sigma\rangle$, where $\langle\sigma\rangle$ is the average diameter of the packing which our unit length. In this work we use packings with sizes distribution shown in Fig. 4, which have polidispersity $19\%$ and $23\%$ for $d=2$ and $d=3$ respectively.
Figure 4: Initial radius distribution for the (a) 2D system and (b) 3D system.
Appendix B The effect of the continuous swap
In order to characterize the dynamics of the system for $d=3$ we consider the usual self-scattering correlation function $F_{s}(k,t)=\left\langle\frac{1}{N}\sum_{i}e^{i\mathbf{k}\cdotp[\mathbf{r}_{j}(t)-\mathbf{r}_{j}(0)]}\right\rangle$, where $\mathbf{r}_{j}(t)$ is the position of the particle $j$ at time $t$ and the wave vector $\mathbf{k}$ satisfies $|\mathbf{k}|=2\pi/\langle\sigma\rangle$. For $d=2$, this definition is not well-suited (long-wavelength vibrational modes bring $F_{s}(k,t)$ to zero for large $t$ and $N$, even in a crystal). This problem is fixed as is usually done by considering the relative motion of particle with respect to their neighbors. It can be achieved by introducing the correlation $C(t)=\frac{1}{N}\sum_{i}C^{i}(t)$, with $C^{i}(t)=\frac{1}{n_{i}(t)}\sum_{\langle ij\rangle}W(1-\frac{|(\mathbf{r}_{j}(t)-\mathbf{r}_{j}(t_{0}))-(\mathbf{r}_{i}(t)-\mathbf{r}_{i}(t_{0}))|}{\langle R\rangle})$ where $\langle R\rangle$ is the average radius of the particles, $n_{i}(t)$ is the number of neighbors of the particle $i$ at time $t$- defined as all particles $j$ for which $|\mathbf{r}_{j}(t)-\mathbf{r}_{i}(t)|<2.4\langle R\rangle$, an estimate of the first shell of neighbors, and $W(x)$ is the Heaviside function. To extract a relaxation time $\tau_{\alpha}$, $F_{s}(k,t)$ and $C(t)$ are fitted with an stretched exponential function $f(t)=\exp(-(t/\tau_{\alpha})^{\beta})$, where $\tau_{\alpha}$ is the relaxation time. The glass packing fraction $\phi_{G}(f_{R},f_{P})$ is then defined such that $\tau_{\alpha}=\tau_{\alpha}^{*}\equiv 10^{7}$ Monte Carlo steps per particle. The associated stretch exponents $\beta_{G}$ are reported in the Appendix as a function of $(f_{R},f_{P})$. Speed up: For $d=3$, our choice of polydispersity and Monte-Carlo algorithm follows closely previous works Berthier et al. (2016); Ninarello et al. (2017), where it was shown that swap can speed up the dynamics by 15 decades or more. Here we observe a giant speed up in our $d=2$ system as well, that continuously builds up as $f_{R}$ decreases toward the swap case $f_{R}=0$ starting from the normal dynamics $f_{R}=1$. Fig.5, in Appendix B, shows the relaxation times $\tau_{\alpha}$ – extracted from correlation functions as recalled in Appendix C – as a function of the packing fraction $\phi$ for different values of $f_{R}$. It is notable that $\tau_{\alpha}$ depends very significantly on $f_{R}$, but that this dependence is continuous. Finally, the speed-up of swap can be approximately extrapolated to the case where the normal dynamics would reach experimental time scales (i.e. would increase by 15 decades). The range of the corresponding $\phi_{G}^{exp}$ can be estimated by fitting the curve $\tau_{\alpha}(\phi)$ of the normal dynamics $f_{R}=1,f_{P}=0$ with plausible functional forms for $\tau_{\alpha}(\phi)$. The functional forms used to fit $\tau_{\alpha}(\phi)$ and estimate $\phi_{G}^{exp}$ are the Vogel-Fulcher-Tamman $\tau_{\alpha}\sim\exp\left(\frac{A}{\phi_{VFT}-\phi}\right)$ or a form with non-singular activation energy $\tau_{\alpha}=\tau^{\prime}_{\infty}\exp\left(A^{\prime}(\phi_{c}-\phi)^{2}\right)$. In agreement with previous such inferences Berthier et al. (2016); Ninarello et al. (2017), we obtain that the swap dynamics $f_{R}=1$ has only slowed-down by 3 to 6 decades at $\phi_{G}^{exp}$: the speed up conferred by swap is very high.
Figure 5: $\tau_{\alpha}/\tau_{\alpha}^{*}$ as a function of $\phi$ for a two-dimensional system for different values of $f_{R}$. The dotted lines shows the values of $\phi_{VFT}$ and $\phi_{P}$ obtained by fitting the normal dynamics simulation ($f_{R}=1$). These fits are shown the insert panel.
Appendix C Stretched exponents
Figures 6-(a,b) show $\beta_{G}$ in color as a function of $f_{R}$ and $f_{P}$ for 2D and 3D system respectively where $\beta_{G}$ is the value of $\beta$ obtained at $\phi_{G}$. These values were obtained through the fitting of function $F_{s}(k,t)$ and $C(t)$ with an stretched exponential function $f(t)=exp(-(t/\tau_{\alpha})^{\beta})$, where $\tau_{\alpha}$ is the relaxation time. In both cases $\beta_{G}$ decreases as the system becomes more restricted.
Figure 6: $\beta_{G}$ as a function of $f_{R}$ and $f_{P}$ for $d=2$ (a) and $d=3$ (b). The color-code represents the value of $\beta_{G}$.
Appendix D Measure of $l_{c}(t)$
To define a coarsening length $l_{c}(t)$, we first define relaxed particles. Particle $i$ is a relaxed particle if $C^{i}(t)\leq 0.5$. Recall that $C^{i}(t)$ represents the proportion of particles $j$ that remain neighbors of particle $i$ after a time $t$. Then we consider that two relaxed particles $i$ and $j$ belong to the same cluster if $|\mathbf{r}_{j}(t)-\mathbf{r}_{i}(t)|<2.4\langle R\rangle$. An example of a growing length scale is presented in Fig.(7) where relaxed particles are shown in red.
Figure 7: Spatio-temporal evolution of the structural relaxation in a 2d, $f_{R}=0,f_{P}=0$ and $\phi=0.79$, for which $\tau_{\alpha}=2.6\times 10^{5}$. Red color indicates relaxed particles. From left to right, the snapshots are taken at times $\tau_{\alpha}/10$ ($l_{c}\approx 2.7$), $\tau_{\alpha}/5$ ($l_{c}\approx 4.5$) and $\tau_{\alpha}/4$ ($l_{c}\approx 5.7$).